OpenStudy (anonymous):

f(x) = x^3 - x Find f'(x) using INCREMENT FORM of the DEFINITION of DERIVATIVE. Thanks.

OpenStudy (anonymous):

f'(x)=[\lim_{h \rightarrow 0}\](f'(x+h)-f(x))/h

OpenStudy (anonymous):

yeah okay see it through guys

OpenStudy (anonymous):

i got 3x^2 + 3xh + h^2

OpenStudy (anonymous):

my equation editor is no working. .

OpenStudy (anonymous):

seems that way

OpenStudy (anonymous):

shut up kiid i need a chain of equalities

OpenStudy (anonymous):

Calyne, now put h=0 in ur final step

OpenStudy (anonymous):

okay so it's 3x^2?

OpenStudy (anonymous):

but isn't d/dx (x^3 - x) = 3x^2 - 1 ..?

OpenStudy (anonymous):

I think you should calculate from the limit step again.

OpenStudy (anonymous):

no, bro, i don't think so.. just fluttering show me

OpenStudy (anonymous):

(x+h)^3-(x+h)-x^3+x x^3+3x^2h+3xh^2+h^3-x^3-x-h+x 3x^2+3xh+h^2-1(Since in the denominator there is a h) 3x^2-1

OpenStudy (rogue):

It should be \[f'(x) = \lim_{h \rightarrow 0} \frac {(x^3 + h)^3 - (x+h) - (x^3 - x)}{h}\]\[f'(x) = \lim_{h \rightarrow 0} \frac {-h + h^3 + 3h^2x + 3hx^2}{h}\]\[f'(x) = \lim_{h \rightarrow 0} = h^2 + 3hx + 3x^2 - 1 = 3x^2 - 1\]

OpenStudy (rogue):

Disregard the equal sign after the lim in the last... fail typing.